OpenStudy (anonymous):

the senior class is taking a trip to an amusement park. They received a special deal where for every 3 tickets they purchased they received one free ticket. 3 tickets cost $53.25. The total purchase of tickets cost; $1384.50. How many tickets did they receive? the senior class is taking a trip to an amusement park. They received a special deal where for every 3 tickets they purchased they received one free ticket. 3 tickets cost $53.25. The total purchase of tickets cost; $1384.50. How many tickets did they receive? @Mathematics

OpenStudy (anonymous):

so first, 1384.50/53.25 = 26. So you have 26 groups of 3 ticket, which means you have 78 tickets. Now you know that for every 3 tickets you buy you get one free. Since you bought 26 groups of 3 tickets, you got 26 additional free tickets. Do total tickets = 78+26=104

OpenStudy (anonymous):

but do you know how to put it in a proportion? :)

OpenStudy (anonymous):

since you know that you paid 1384.50 total for all your tickets, and a group of 3 tickets costs 53.25. Divide the total amount spent by the cost of a group of 3 tickets. Then you know how many groups of three tickets you have, which is how many additional free tickets you receive.

OpenStudy (anonymous):